OpenStudy (anonymous):

lim t->infinity of (log(t-3)-log(t^2+6t+9))

OpenStudy (anonymous):

I think i get indeterminant form infinity-infinity, so i make it log(t-3)/(-1/log(t^2+6t+9)??

OpenStudy (anonymous):

no use property of logs

OpenStudy (anonymous):

rewrite as \[\log(\frac{t-3}{t^2+6t+9})\]

OpenStudy (slaaibak):

negative infinity.

OpenStudy (anonymous):

Would going the other way, as I mentioned, work? Or is that headed int he wrong direction completely?

OpenStudy (anonymous):

then \[\lim_{t \rightarrow \infty}\frac{t-3}{t^2+6t+9}=0\] and so your limit, as sheg said, is \[-\infty\]

OpenStudy (anonymous):

i think you are confusing that with the method for solving \[0\times \infty\] for example where you would rewrite the expression. but in this case you have \[\infty - \infty\] and that is different

OpenStudy (anonymous):

I'm quite certain infinity-infinity is an indeterminant form? Okay so doing your method, how did you get rid of the log? and why did you set it equal to 0?

OpenStudy (anonymous):

\[\log((t-3)/(t^2+6t+9)=\log((t-3)/(t+3)(t+3))\] But how can i simplify that?